# Protocol Futuring: Speculating Second-Order Dynamics of Protocols in Sociotechnical Infrastructural Futures

### v1

@@ Abstract.
Drawing on infrastructure studies in HCI and CSCW, this paper introduces Protocol Futuring, a methodological framework that extends design futuring by foregrounding protocols—rules, standards, and coordination mechanisms—as the primary material of speculative inquiry. Rather than imagining discrete future artifacts, Protocol Futuring examines how protocol rules accumulate drift, jam, and other second-order effects over long temporal horizons. We demonstrate the method through a case study of Knowledge Futurama, a multi-team participatory workshop exploring millennial-scale knowledge preservation. Using a relay format in which teams inherited and reinterpreted partially formed designs, the workshop revealed how ambiguous handovers, adversarial reinterpretations, shifting cultural norms, and crisis dynamics transform protocols as they move across communities and epochs. The case shows how Protocol Futuring makes infrastructural politics and long-run consequences analytically visible. We discuss the method’s strengths, limitations, and implications for researchers seeking to investigate emergent sociotechnical systems whose impacts unfold over extended timescales.
Protocol Design, Protocol Studies, Sociotechnical Systems, Infrastructure Studies, Design Futuring, Research Through Design, Speculative Design, Critical Computing
$c$s: Human-centered computing Collaborative and social computing theory, concepts and paradigms
@@ 1. Introduction
In a very real sense, “the fate of our world may rest on protocols”(Bennke2025Protological). Technical protocols—the standards and rule sets that orchestrate reliable interactions among AI agents—underpin planetary-scale computation and coordination (Hui2024Machine; Bratton2015stack). They range from floating-point arithmetic used by every computing device (Lee2025Floating), to communication protocols such as HTTP and TCP/IP, and cryptographic primitives and smart contracts like Ethereum (Zouari2025Ethereum). Unlike products or platforms, protocols tend to “disappear” into infrastructure (Star1996Steps), their influence surfacing only when failures or side-effects emerge. This raises a design challenge: how can we proactively envision the futures shaped by protocols, not just the applications built atop them?
@@
Over the last two decades, HCI researchers have widely adopted the method of ”design futuring” (Celik2025Reviewing; fry2009design; Bardzell2012Critical), such as speculative design (Auger2013Speculativea) and design fiction (Blythe2014Research) as tools to create conceptual artifacts and narratives to probe alternative sociotechnical worlds (dunne2013speculative). Critical and speculative design (Bardzell2012Critical) projects seek to probe different social values, norms, and material configurations. Yet most remain artifact-centric: they foreground near-term devices and user interactions (bleecker2009design), while the “invisible” infrastructures of rules and standards stay in the background (wong2020infrastructural). Long-term issues of evolution, maintenance, and governance thus remain implicit, prompting calls for “infrastructural speculation” (wong2020infrastructural). This gap has led to calls for infrastructural speculation in design research. Wong et al. (wong2020infrastructural) ask how speculative design can “explicitly center and raise questions about the infrastructures within which products are entangled”.
@@
Protocols (summerofprotocols2023) differ from products in that their effects are usually indirect and emergent. A well-designed protocol fades into the background of a system; people interact through it, not with it. Its implications unfold through second-order dynamics in the sociotechnical system, rather than through one-to-one user interactions. For instance, when a new data-sharing protocol is widely adopted, it might initially improve interoperability (a first-order intended effect), but it can also lead to second-order phenomena: drift, where stakeholders adapt or game the system (as encapsulated in Goodhart’s law: “[w]hen a measure becomes a target, it ceases to be a good measure” (Manheim2019Categorizing)); jams, where scaling up usage creates congestion or conflicts (echoing Frederik Pohl’s dictum that a good futurist anticipates “not the automobile but the traffic jam” (pohl1973)); or ossification, where an entrenched standard becomes resistant to change, locking in certain behaviors or power structures. These emergent effects often only become apparent when protocols interact with complex social contexts – users, institutions, and other technologies that repurpose or stretch the rules beyond their original intent. Traditional user-centered design methods struggle to surface such long-term, systemic implications. This is why protocol futuring is necessary to systematically explore and speculate about these hidden dynamics before they become tomorrow’s crises or lock-ins.
@@
In this paper, we introduce Protocol Futuring as a methodological framework for critical computing and design research for infrastructure studies. Protocol Futuring shifts speculative inquiry away from singular gadgets and toward the infrastructural rules and standards that coordinate sociotechnical futures. It is a framework rather than a single method: it can be realized through design fiction, participatory workshops, AI-based simulation, adversaries contests, and other research‑through‑design techniques. We define Protocol Futuring through five core commitments:
@@
  • Second‑order dynamics: probe drift, jams, feedbacks, and ossification—not only intended use.
  • Situated lifeworlds: embed protocols in social, organizational, legal, and environmental contexts rather than abstract deployments (wong2020infrastructural).
  • Multi‑stakeholder argumentation: surface protocols as negotiated settlements that encode values and distribute benefits/risks.
  • Research‑through‑Design & reflexivity: think‑by‑making protocol probes, iteratively reflect on methods and positionality.
In operationalizing these principles, we provide a workflow that guides researchers through the process of Protocol Futuring. identifying a focal protocol or infrastructural question; constructing speculative scenarios or prototypes around that protocol; engaging stakeholders or participants to inhabit and stress-test those scenarios; and reflecting on the emergent dynamics and insights gained. We detail this process in Section 3, offering practical steps and considerations for applying Protocol Futuring in design research projects.
@@
In summary, this paper contributes Protocol Futuring as a novel methodological framework for HCI, CSCW, and design research. We define the concept of Protocol Futuring and demonstrate its application across multiple scenarios, providing an infrastructural lens for speculating about sociotechnical futures. Our work offers design researchers a framework (with concrete principles and tactics) for thinking beyond individual gadgets to the rules and institutions that drive technology’s long-term impacts. By focusing on protocols and their second-order dynamics, Protocol Futuring reveals hidden politics and complex entanglements that might otherwise remain invisible in product-centered futuring.
We argue that this approach enriches critical computing practice by allowing researchers to engage with questions of governance, maintenance, and systemic change in envisioned worlds. Finally, we reflect on the strengths of Protocol Futuring – such as its power to surface infrastructural power imbalances – as well as its challenges (e.g. how to validate or assess speculative insights about systems that do not yet exist).
@@ 2. Background
We ground the development of protocol futuring in three strands of background literature. First, we survey the emerging field of protocol studies, which examines protocols as sociotechnical phenomena and argues for their growing importance. This informs our understanding of what “protocol” means beyond the purely technical definition, and why protocols warrant speculative exploration. Second, we review work in speculative design, design fiction, and participatory design, tracing how these practices have evolved from focusing on artifact-centric futures towards engaging with infrastructure, policy, and plural perspectives. This situates our method among efforts to make design fictions more systemic and inclusive. Third, we draw on infrastructure studies in HCI and STS, which provide concepts (e.g. infrastructural inversion, infrastructuring, and themes of growing/appropriating/coping) that parallel our focus on the long-term life of technologies. Integrating these threads, we articulate how protocol futuring can serve as a form of critical infrastructural speculation oriented to both technical and social dimensions of future networks.
@@ 2.1.1. Technical Protocol Design Practices
Technical protocols are the backbone of interoperable technology, prescribing how systems communicate and function. Classic Internet design principles exemplify this culture of protocol development. The Internet Engineering Task Force (IETF) famously embraces “rough consensus and running code,” a maxim that privileges practical implementations and community agreement over formal authority (clark1992tussle). David Clark’s dictum—“We reject: kings, presidents, and voting. We believe in: rough consensus and running code”—encapsulates the preference for open, collaborative iteration in protocol design. Protocols such as TCP/IP were deliberately minimal, “caring nothing for what is said or who said it,” to remain flexible building blocks (saltzer1984end). These values of simplicity, modularity, and extensibility have been credited with the Internet’s success, even as they separated social or policy considerations from early engineering work (denardis2014global).
2.1.2. Social and Governance Protocols
@@
Computing definitions treat protocols as rules for data transmission, but interdisciplinary scholarship broadens the concept to sociotechnical coordination mechanisms—“codified behaviors” enabling complex, distributed activity (galloway2004protocol). The Summer of Protocols initiative (2023) adopted and debated Danny Ryan’s definition of a protocol as “a stratum of codified behavior that allows for the construction or emergence of complex coordinated behaviors at adjacent loci,” eventually condensing it to the pithy insight that “a protocol is an engineered argument” (summerofprotocols2023). In this view, every protocol encodes a negotiation of values and priorities, shaping collaboration and conflict alike.
@@
Protocols often fade into the background when functioning well, a phenomenon Alfred North Whitehead anticipated when he wrote that “civilization advances by extending the number of important operations we can perform without thinking of them”(whitehead1911). They coordinate actions at scales ranging from the mundane (traffic lights, handshakes) to the global (TCP/IP, blockchain consensus), yet typically draw attention only when they fail(star1999ethnography). Scholars stress that protocols are infrastructural forces on par with markets or organizations but remain undertheorized and under-taught (pipek2009infrastructuring). Engineers may not learn to read or write protocol specifications formally, and most social-science frameworks do not treat protocols as first-class analytical objects.
@@
Despite their hiddenness, protocols carry social and political weight. Lawrence Lessig’s dictum that “code is law” captures how software rules can regulate behavior as effectively as legal statutes (lessig1999code). Contemporary analyses extend this logic to “social protocols”—the explicit and tacit rules that structure interaction online and off (warden2023social). Platform algorithms, moderation policies, and community norms jointly govern information flows, demonstrating how technical and social layers of protocol interlock. Formal standard-making bodies increasingly treat social values as matters of codification: ISO 14001 and ISO 26000, for example, establish environmental and social responsibility protocols beyond conventional law (Yates2019).
@@
Yet the very success of protocols breeds neglect. Effective protocols recede into infrastructure, making failures hard to anticipate and reforms difficult to introduce. Internet history offers cautionary examples of “protocol ossification,” where widely deployed standards become so entrenched that even beneficial upgrades face rejection by middleboxes or firewalls (handley2017ossification). More generally, neglected protocols can grow fragile, sclerotic, or vulnerable to capture by narrow interests or malicious actors (asparouhova2023protocols). Conversely, well-designed and well-stewarded protocols can enable major leaps in coordination, as rail gauges, electrical standards, and Internet protocols did in earlier eras.
@@ 2.1.3. Standardization Processes: History and Institutional Context
Protocols and standards emerge through institutionalized negotiations that balance innovation, consensus, and power. Yates and Murphy trace three historical “waves” of global standard setting (Yates2019). In the late 19th and early 20th centuries, professional societies such as ASTM and IEEE developed technical specifications to enable industrial interoperability. Mid-20th-century efforts like the ISO and ITU expanded coordination globally. From the late 1990s onward, a third wave has addressed broad governance issues—human rights, environmental stewardship, social responsibility—treating values themselves as subjects of standardization.
@@
Different arenas exemplify different governance models. The IETF relies on an open, volunteer-driven RFC process, while ISO and ITU operate formal, consensus-based negotiations among governments and corporations. Each approach trades off agility and political legitimacy. The enduring challenge, as Hanseth et al. noted, is balancing stability and flexibility so that standards remain universal yet adaptable (hanseth1996tension). Rigid standards can become barriers to change, while excessive flexibility undermines interoperability.
@@
Standards can also spur creativity rather than stifle it. Jackson and Barbrow’s study of an ecology research network showed that defining shared data and workflow protocols was a generative design activity, coining the idea of “standards and/as innovation” (jackson2015standards). By bringing diverse actors together to formalize solutions, standardization can itself be a site of discovery and collaboration.
2.1.4. Speculative and Fictional Protocols in Design and HCI
Design fiction and speculative design extend protocol thinking into future-oriented inquiry. By creating fictional scenarios, artifacts, or narratives, these approaches reveal latent dependencies and open space for imagining alternatives. Lyu et al. (2025) survey how HCI researchers have used speculation to probe future infrastructures (lyu2025infrastructure). Examples include Rodden et al. (2013) using “agents” to simulate smart energy grids (rodden2013smart), Steinhardt and Jackson (2015) analyzing “anticipation work” in ecology and ocean science (steinhardt2015anticipation), Soden and Kauffman (2019) on “infrastructuring the imaginary”(soden2019infrastructuring), and Jabbar and Bjørn (2019) examining how blockchain reconfigures shipping logistics(jabbar2019blockchain).
Other projects use absence as a provocation: Grandhi et al. (2020) asked “An Internet-less World?” to elicit reflection on society’s dependence on network protocols (grandhi2020internetless). In the blockchain space, workshops like the Ethereum Foundation’s “protocol futuring” explore how tweaking basic consensus or economic parameters might unleash unforeseen social and technical dynamics. Composable Life, for instance, imagines blockchain as an “unstoppable nature,” using speculative narrative to interrogate long-run protocol evolution (hu2024composable).

### v2

@@ Abstract.
Drawing on infrastructure studies in HCI and CSCW, this paper introduces Protocol Futuring, a methodological framework that extends design futuring by foregrounding protocols—rules, standards, and coordination mechanisms—as the primary material of speculative inquiry. Rather than imagining discrete future artifacts, Protocol Futuring examines how protocol rules accumulate drift, jam, and other second-order effects over long temporal horizons. We demonstrate the method through a case study of Knowledge Futurama, a multi-team participatory workshop exploring millennial-scale knowledge preservation. Using a relay format in which teams inherited and reinterpreted partially formed designs, the workshop revealed how ambiguous handovers, adversarial reinterpretations, shifting cultural norms, and crisis dynamics transform protocols as they move across communities and epochs. The case shows how Protocol Futuring makes infrastructural politics and long-run consequences analytically visible. We discuss the method’s strengths, limitations, and implications for researchers investigating emergent sociotechnical systems whose impacts unfold over extended timescales.
Protocol Design, Protocol Studies, Sociotechnical Systems, Infrastructure Studies, Design Futuring, Research Through Design, Speculative Design, Critical Computing, Drift, Second-Order Dynamics
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790527isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Collaborative and social computing design and evaluation methods
Figure 1. Protocol Futuring focuses on protocols as the primary material of speculative inquiry and operationalizes long-term infrastructural speculation through iterative handoffs between teams. In the Knowledge Futurama workshop, participants alternated between “Blue Build” phases (designing preservation protocols) and “Red Crisis” phases (introducing disruptions), passing partially-formed designs to new teams across simulated centuries. This relay structure makes visible how protocols drift, jam, and ossify as they traverse multiple hands with incomplete context—surfacing second-order dynamics that artifact-centric design futures often miss.
Protocol Futuring focuses on protocols as the primary material of speculative inquiry and operationalizes long-term infrastructural speculation through iterative handoffs between teams. In the Knowledge Futurama workshop, participants alternated between “Blue Build” phases (designing preservation protocols) and “Red Crisis” phases (introducing disruptions), passing partially-formed designs to new teams across simulated centuries. This relay structure makes visible how protocols drift, jam, and ossify as they traverse multiple hands with incomplete context—surfacing second-order dynamics that artifact-centric design futures often miss.
@@ 1. Introduction
In a very real sense, “the fate of our world may rest on protocols” (Bennke and Schneider, 2025). Technical protocols—the standards and rule sets that orchestrate reliable interactions among humans and agents—underpin planetary-scale computation and coordination (Hui, 2024; Bratton, 2016). They range from floating-point arithmetic used by every computing device (Lee and Hyland, 2025), to communication protocols such as HTTP and TCP/IP, and cryptographic primitives and smart contracts like Ethereum (Zouari et al., 2025). Unlike products or platforms, protocols tend to “disappear” into infrastructure (Star and Ruhleder, 1996), their influence surfacing only when failures or side-effects emerge. This raises a design challenge: how can we proactively envision the futures shaped by protocols, not just the applications built atop them?
@@
Over the last two decades, HCI researchers have widely adopted the method of “design futuring” (Celik and Kaya, 2025; Fry, 2009; Bardzell et al., 2012), such as speculative design (Auger, 2013) and design fiction (Blythe, 2014) as tools to create conceptual artifacts and narratives to probe alternative sociotechnical worlds (Dunne and Raby, 2013). Critical and speculative design (Bardzell et al., 2012) projects seek to probe different social values, norms, and material configurations. Yet most remain artifact-centric: they foreground near-term devices and user interactions (Bleecker, 2009), while the “invisible” infrastructures of rules and standards stay in the background. Long-term issues of evolution, maintenance, and governance thus remain implicit, prompting calls for “infrastructural speculation” (Wong et al., 2020): Wong et al. (2020) ask how speculative design can “explicitly center and raise questions about the infrastructures within which products are entangled”.
@@
Protocols (Summer of Protocols, 2025) differ from products in that their effects are usually indirect and emergent. A well-designed protocol fades into the background of a system; people interact through it, not with it. Its implications unfold through second-order dynamics in the sociotechnical system, rather than through one-to-one user interactions. For instance, when a new data-sharing protocol is widely adopted, it might initially improve interoperability (a first-order intended effect), but it can also lead to second-order phenomena: drift, where stakeholders adapt or game the system, as encapsulated in Goodhart’s law: “when a measure becomes a target, it ceases to be a good measure” (Goodhart, 1984; Manheim and Garrabrant, 2018); jams, where scaling up usage creates congestion or conflicts, echoing Frederik Pohl’s observation that “a good science fiction story should be able to predict not the automobile but the traffic jam” (Pohl, 1968); or ossification, where an entrenched standard becomes resistant to change, locking in certain behaviors or power structures (Ammar, 2018). These emergent effects often only become apparent when protocols interact with complex social contexts—users, institutions, and other technologies that repurpose or stretch the rules beyond their original intent. Traditional user-centered design methods struggle to surface such long-term, systemic implications.
@@
To systematically explore and speculate about these hidden dynamics before they become tomorrow’s crises or lock-ins, in this paper, we introduce “Protocol Futuring” (PF) as a methodological framework for critical computing and design research for infrastructure studies. Protocol futuring shifts speculative inquiry away from singular gadgets and toward the infrastructural rules and standards that coordinate sociotechnical futures. It is a framework rather than a single method: it can be realized through design fiction, participatory workshops, AI-based simulation, adversarial contests, and other research-through-design techniques (Blythe, 2014). We define Protocol Futuring through five core commitments:
@@
  • Second-order dynamics: probe drift, jams, and ossification beyond intended use.
  • Situated lifeworlds: embed protocols in social, organizational, legal, and environmental contexts rather than abstract deployments.
  • Multi-stakeholder argumentation: surface protocols as negotiated settlements that encode values and distribute benefits/risks.
  • Research-through-Design and Reflexivity: think-by-making protocol probes, iteratively reflect on methods and positionality.
In operationalizing these principles, we provide a workflow that guides researchers through the process of Protocol Futuring: identifying a focal tension or infrastructural problem; constructing speculative scenarios or prototypes around that protocol to ease the tension or solve the problem; engaging stakeholders or participants to inhabit and stress-test those scenarios; and reflecting on the emergent second-order dynamics and insights gained. We detail this process in Section 3, offering practical steps and considerations for applying Protocol Futuring in design research projects.
@@
In summary, this paper contributes Protocol Futuring as a novel methodological framework for HCI, CSCW, and design research. We define the concept of Protocol Futuring and demonstrate its application across multiple scenarios, providing an infrastructural lens for speculating about sociotechnical futures. Our work offers design researchers a framework (with concrete principles and tactics) for thinking beyond individual gadgets to the rules and institutions that drive technology’s long-term impacts. By focusing on protocols and their second-order dynamics, Protocol Futuring reveals hidden politics and complex entanglements that might otherwise remain invisible in product-centered futuring. We argue that this approach enriches critical computing practice by allowing researchers to engage with questions of governance, maintenance, and systemic change in envisioned worlds. Finally, we reflect on the strengths of Protocol Futuring—such as its power to surface infrastructural power imbalances—as well as its challenges, e.g. how to validate or assess speculative insights about systems that do not yet exist.
@@ 2. Background
We ground the development of protocol futuring in three strands of background literature. First, we survey the emerging field of protocol studies, which examines protocols as sociotechnical phenomena and argues for their growing importance. This informs our understanding of what “protocol” means beyond the purely technical definition, and why protocols warrant speculative exploration. Second, we review work in speculative design, design fiction, and participatory design, tracing how these practices have evolved from focusing on artifact-centric futures towards engaging with infrastructure, policy, and plural perspectives. This situates our method among efforts to make design fictions more systemic and inclusive. Third, we draw on infrastructure studies in HCI and STS, which provide concepts that parallel our focus on the long-term life of technologies. Integrating these threads, we articulate how protocol futuring can serve as a form of critical infrastructural speculation oriented to both technical and social dimensions of future networks.
@@ 2.1.1. Technical Protocol Design Practices
Technical protocols are the backbone of interoperable technology, prescribing how systems communicate and function. Classic Internet design principles exemplify this culture of protocol development. The Internet Engineering Task Force (IETF) famously embraces “rough consensus and running code,” a maxim that privileges practical implementations and community agreement over formal authority (Russell, 2006). David Clark’s dictum—“We reject: kings, presidents, and voting. We believe in: rough consensus and running code”—encapsulates the preference for open, collaborative iteration in protocol design. Protocols such as TCP/IP were deliberately minimal, “caring nothing for what is said or who said it,” to remain flexible building blocks (Galloway, 2004). These values of simplicity, modularity, and extensibility have been credited with the Internet’s success, even as they separated social or policy considerations from early engineering work (DeNardis, 2014).
2.1.2. Protocols as Sociotechnical Infrastructures
@@
Computing definitions treat protocols as rules for data transmission, but interdisciplinary scholarship broadens the concept to sociotechnical coordination mechanisms—“codified behaviors” enabling complex, distributed activity (Galloway, 2004). The Summer of Protocols initiative [^1] https://summerofprotocols.com/ adopted and debated Danny Ryan’s definition of a protocol as “a stratum of codified behavior that allows for the construction or emergence of complex coordinated behaviors at adjacent loci,” eventually condensing it to the pithy insight that “a protocol is an engineered argument” (Summer of Protocols, 2025). In this view, every protocol encodes a negotiation of values and priorities, shaping collaboration and conflict alike.
@@
Protocols often fade into the background when functioning well, a phenomenon Alfred North Whitehead anticipated when he wrote that “civilization advances by extending the number of important operations we can perform without thinking of them” (Whitehead, 1911). They coordinate actions at scales ranging from the mundane (traffic lights, handshakes) to the global (TCP/IP, blockchain consensus), yet typically draw attention only when they fail (Star, 1999). Scholars stress that protocols are infrastructural forces on par with markets or organizations but remain undertheorized and under-taught (Pipek and Wulf, 2009). Engineers may not learn to read or write protocol specifications formally, and most social-science frameworks do not treat protocols as first-class analytical objects.
@@
Despite their hiddenness, protocols carry social and political weight. Lessig (1999)’s dictum that “code is law” captures how software rules can regulate behavior as effectively as legal statutes. Contemporary analyses extend this logic to “social protocols”—the explicit and tacit rules that structure interaction online and off (Warden, 2023). Platform algorithms, moderation policies, and community norms jointly govern information flows, demonstrating how technical and social layers of protocol interlock. Masnick (2019) argues that shifting from proprietary platforms to open protocols could address persistent challenges in content moderation and free speech online, proposing that protocol-based architectures give users more choice and create healthier competitive dynamics than centralized platform governance. Formal standard-making bodies increasingly treat social values as matters of codification. ISO 14001 (“Environmental management systems”) and ISO 26000 (“Social responsibility”), for example, establish environmental and social responsibility protocols that go beyond conventional law (Yates and Murphy, 2019).
@@
Yet the very success of protocols breeds neglect. Effective protocols recede into infrastructure, making failures hard to anticipate and reforms difficult to introduce. Internet history offers cautionary examples of “protocol ossification,” (Ammar, 2018) where widely deployed standards become so entrenched that even beneficial upgrades face rejection by middleboxes or firewalls (Goldberg, 2014). More generally, neglected protocols can grow fragile, sclerotic, or vulnerable to capture by narrow interests or malicious actors. Askonas (2025) describes how protocols create “accountability sinks”—situations where decisions are delegated to policies or automated systems such that no human appears responsible or in charge. In a protocol-governed society, agency becomes invisible: the more decentralized and scalable the system, the more responsibility evaporates, even as power continues to operate (Masnick, 2019). Conversely, well-designed and well-stewarded protocols can enable major leaps in coordination, as rail gauges, electrical standards, and Internet protocols did in earlier eras.
@@ 2.1.3. Standardization Processes: History and Institutional Context
Protocols and standards emerge through institutionalized negotiations that balance innovation, consensus, and power. Yates and Murphy (2019) trace three historical “waves” of global standard setting. In the late 19th and early 20th centuries, professional societies such as ASTM and IEEE developed technical specifications to enable industrial interoperability. Mid-20th-century efforts like the ISO and ITU expanded coordination globally. From the late 1990s onward, a third wave has addressed broad governance issues—human rights, environmental stewardship, social responsibility—treating values themselves as subjects of standardization.
@@
Different arenas exemplify different governance models. The IETF relies on an open, volunteer-driven RFC process, while ISO and ITU operate formal, consensus-based negotiations among governments and corporations. Each approach trades off agility and political legitimacy. The enduring challenge, as Hanseth et al. (1996) noted, is balancing stability and flexibility so that standards remain universal yet adaptable. Rigid standards can become barriers to change, while excessive flexibility undermines interoperability.
@@
Standards can also spur creativity rather than stifle it. Jackson and Barbrow (2015)’s study of an ecology research network showed that defining shared data and workflow protocols was a generative design activity, coining the idea of “standards and/as innovation”. By bringing diverse actors together to formalize solutions, standardization can itself be a site of discovery and collaboration.
2.1.4. Speculative and Fictional Protocols
Recently, Protocol fiction has emerged as a distinct approach that foregrounds protocols themselves as the primary object of speculation. Webb (2023) defines protocol fiction as a method for imagining new infrastructure—such as drone delivery networks or health screening ecosystems—by specifying protocols that enable cooperation among aligned actors and articulating both the belief (a plausible path to implementation) and desire (a compelling visualization of the future) needed for adoption. Unlike conventional design fiction, which often centers on discrete artifacts, protocol fiction treats the coordination mechanism itself as the speculative material.
This orientation toward protocols as narrative material has gained traction through initiatives like the Terminological Twists fiction contest, which challenged writers to take existing protocol-related technical concepts and reverse-engineer them into compelling science fiction stories (Protocolized, 2025b). The contest, which attracted fifty submissions, exemplifies protocol fiction as “a fresh subgenre of sci-fi, with systems and rules almost as protagonists in their own right” (Protocolized, 2025a). Guided by what practitioners call “Chiang’s Law”—that science fiction is about strange rules while fantasy is about special people—protocol fiction explores regimes of rules and the dynamics of worlds that grow on those protocols. Winning entries such as “Noise Ordinance,” “The 40-Hour Work Week,” and “DHCP” demonstrate how technical protocol concepts can anchor speculative narratives that probe social and political implications.
The Summer of Protocols initiative employs protocol fiction as a means of provocation and exploration. For example, Composable Life envisions blockchain as an “unstoppable nature” that gives rise to artificial life, leveraging speculative narrative and immersive virtual reality to probe the long-term evolution of protocols and the intricate, symbiotic relationships between humans, AI, and blockchain technology (Hu and Fangting, 2025). In a similar vein, Autonomous Realities introduces speculative protocols for shared mixed reality experiences, including an embodied protocol for establishing MR sessions inspired by ancient secret handshakes, alongside a meta-protocol for sustaining persistent, autonomous shared reality layers (Hu, 2025). These projects push protocol fiction beyond traditional narrative, extending it into the realms of experiential and embodied interaction.
HCI research has actively engaged with speculative approaches to protocols and infrastructure. For instance, Rodden et al. (2013) deployed “agents” within homes to simulate future smart energy grids, probing how people might interact with and perceive emerging energy infrastructures. Soden and Kauffman (2019) explored how communities practice “infrastructuring the imaginary” in response to challenges like sea-level rise, highlighting the imaginative work underpinning collective adaptation. Similarly, Jabbar and Bjørn (2019) demonstrated that introducing blockchain technologies compels organizations to reimagine workflows and reshape their infrastructural imaginaries. Some projects take an inverse approach, using the absence of protocols as a critical provocation: Grandhi et al. (2020) posed the question “An Internet-less World?” to prompt reflection on society’s profound dependency on network protocols, illustrating how scenarios of protocol failure can serve as productive sites for creative inquiry.